# Mutual Information Surprise: Rethinking Unexpectedness in Autonomous Systems

### v1

@@ 5 Conclusion
In this work, we reimagined the concept of surprise as a mechanism for fostering understanding, rather than merely detecting anomalies. Traditional definitions—such as Shannon and Bayesian Surprises—focus on single-instance deviations and belief updates, yet fail to capture whether a system is truly growing in its understanding over time. By introducing Mutual Information Surprise (MIS), we proposed a new framework that reframes surprise as a reflection of learning progression, grounded in mutual information growth.
We developed a formal test sequence to monitor deviations in estimated mutual information, and introduced a reaction policy, MISRP, that transforms surprise into actionable system behavior. Through a synthetic case study and a real-time pollution map estimation task, we demonstrated that MIS governance offers clear advantages over conventional sampling strategies. Our results show improved stability, better responsiveness to environmental drift, and significant reductions in estimation error. These findings affirm MIS as a robust and adaptive supervisory signal for autonomous systems.
@@
$=\frac{n-1}{n}\log(n-1)-\frac{n-2}{n}\log n$
$=\frac{n-1}{n}(\log n+\log(1-\frac{1}{n}))-\frac{n-2}{n}\log n$
$=\frac{n-1}{n}\left[\log n-\frac{1}{n}-\frac{1}{2n^{2}}+\mathcal{O}\left(\frac{1}{n^{3}}\right)\right]-\frac{n-2}{n}\log n$
$=\frac{1}{n}\log n-\frac{1}{n}+\frac{1}{2n^{2}}+\mathcal{O}\left(\frac{1}{n^{3}}\right)$
$=\frac{\log n-1}{n}+\frac{1}{2n^{2}}+\mathcal{O}\left(\frac{1}{n^{3}}\right)\leq\frac{\log n}{n}.$
@@
2:while $i<k$ (ii starts from 0do
@@
13:   while $j\leq t$ ($j$ starts from 0do

### v3

@@ 5 Conclusion
We started the paper by presenting a vivid picture of the recent race towards autonomous systems. We argued that while the bodies of various autonomous system have advanced visibly, even to the general public, over the past decade, the brains of these autonomous systems still miss a critical element, which is how to define and react to surprises. Unlike classical surprise measures that merely characterize statistical irregularities, we re-imagine the concept of surprise as a mechanism for fostering understanding. We further define a new surprise metric based on mutual information and re-frame surprise as a reflection of learning progression grounded in mutual information growth.
We developed a formal test sequence to monitor deviations in the estimated mutual information, and introduced a reaction policy, MISRP, that transforms surprise into actionable system behavior. Through a synthetic case study and a pollution map estimation task, we demonstrated that MIS governance offers clear advantages over conventional sampling strategies. Our results show improved stability, better responsiveness to environmental drift, and significant reductions in estimation error. These findings affirm MIS as a robust and adaptive supervisory signal for autonomous systems.
@@
$\leq\frac{n-1}{n}\log\frac{n}{n}+\frac{1}{n}\log n=\frac{\log n}{n}.$
@@
In the last simulation experiment with the pollution map, we use the same PDE with modified parameters. Specifically, the pollution sources $S(\mathbf{x})$ is removed, and the decay parameter $\zeta$ is reduced to $0.1$ in the second phase.
@@
2:while $i<k$ (ii starts from 0) do
@@
13:   while $j\leq t$ ($j$ starts from 0) do
@@
Non-smoothed Error Progression with Action Overlays
Here we present the non-smoothed estimation error progression figure with action overlays.
Figure 11: A non-smoothed visualization of estimation error progression with MISRP action overlays.